I need help sovling
7-5x=-(x+1)
Thank you!!

Found 2 solutions by jim_thompson5910, ewatrrr:
Answer by jim_thompson5910(35256)

7-5x=-(x+1)


7-5x=-x-1


-5x+x=-1-7


-4x=-8


x=-8/(-4)


x = 2

Answer by ewatrrr(24785)

 
Hi,
Removing parenthesis and isolating x on one side of the Eq to solve:
Question states***
7-5x = -(x+1)
7-5x = -x-1
8 = 4x
2 = x OR x = 2!
CHECKING our Answer***
7 -10 = -(3) = -3